A plate, a dish and a cup are on a table. The plate is square with a side 30 cm long. The dish is 20 cm in diameter, and the cup has a 15 cm diameter. What is the probability of tossing a grape into the cup and not onto the plate or the dish?
13.4 14.6 15.6 12.7
i am doing something wrong because i dont come up with one of the choices.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

A plate, a dish and a cup are on a table. The plate is square with a side 30 cm long. The dish is 20 cm in diameter, and the cup has a 15 cm diameter. What is the probability of tossing a grape into the cup and not onto the plate or the dish?
--------
Area of the square: 30^2 = 900 cm^2
Area of the dish: pi*10^2 = 100pi cm^2
Area of the cup: pi*7.5^2 = 56.25pi cm^2
----------------------------------------
P(cup) = 56.25pi/{900 + 100pi) = 176.7146/(1214.16) = 0.146..
=========================
Cheers,
Stan H.
==============================
13.4 14.6 15.6 12.7

Answer by solver91311(24713)



Calculate the area of each of the three items. Divide the area of the cup by the sum of the three areas.

In fact, none of the answers is correct as stated. Probabilities are ALWAYS in the range . Now if the answers had been given as 13.4% 14.6% 15.6% 12.7%, that would be another story.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism